# references.bib

### v1

@@ 1. Introduction
Support for management and partitioning of shared system resources such as shared cache and memory bandwidth has recently become available in commercial off-the-shelf x86 server processors. The features, which we refer to as hardware-based quality-of-service (QoS) enforcement features, may be dynamically controlled at runtime using processor model-specific registers or an operating system-level interface. They enable systems to enforce availability guarantees and mitigate the interference caused by shared resource congestion. The benefits of intelligently tuning such features have been shown in previous work to include increased system performance (lo_heracles_2015), improved system utilization (patel_clite_2020), and reduced power consumption (nejat_cooperative_2022). However, to the best of our knowledge, there has been no previous attempt at surveying the academic and industrial landscapes to understand the applicability of these new features in different contexts.
@@
One such context of interest is that of general-purpose cloud computing. The cloud-connected nature of day-to-day applications implies that performance issues of cloud-based web services now have a prevalent impact on the daily lives of the general population in a way that was unthinkable a few decades ago. In fact, a (ballard_industry_2023) US-based survey indicated that $48\%$ of companies use industry-tailored clouds with another $20\%$ intending to make the transition to cloud-based solutions within 12 months (ballard_industry_2023). Latency-sensitive web services are often referred to as latency critical (LC) tasks, which is one of two broad categories cloud-based workloads may be classified into. The response times of LC services often directly affect the end-user experience and are therefore monitored and kept to performance targets specified by service level objectives (SLOs). Best-effort (BE) jobs on the other hand are time non-critical, predictable workloads that may run whenever sufficient resources become available, commonly using a batch-format scheduling system. Thus, such jobs are well suited to improve server utilization by consuming resources left unused by LC services based on current loads, a technique leveraged by several previous works (lo_heracles_2015; chen_parties_2019; nikas_dicer_2019; patel_clite_2020; papadakis_improving_2017; qin_locpart_2017; qiu_slo_2021; zhao_rhythm_2020; li_rldrm_2020; zhu_dirigent_2016; chen_avalon_2019). BE tasks such as machine learning model training or big data analysis jobs are often data intensive and consequently disproportionately “noisy” users of shared system resources. Unless mitigated, this congestion may negatively affect the performance of LC services (dasari_identifying_2013) and subsequently, end user experience.
@@
The congestion of shared resources such as last-level cache (LLC) and memory bandwidth (MB) caused by the co-location of workloads on multicore systems has been shown in previous studies to reduce system performance (dasari_identifying_2013; lofwenmark_understanding_2016; zhuravlev_addressing_2010). The importance of this becomes evident when considering the SLOs that web services often are required to meet. Not reaching performance targets due to resource congestion may end up costly as BE jobs may need to be canceled and re-queued. Alternatively, additional computing resources may need to be provisioned if permitted by the underlying infrastructure platform. Hardware-based QoS enforcement features are designed to mitigate the issues of shared resource congestion. However, hardware-based QoS enforcement is indirect in nature (see Section 2) and thus, intelligent tuning is required to achieve predictable and performant service operations in cloud environments.
@@ 1.1. Our Contributions
    We survey a search space of $3685$ articles and identify $79$ papers that fit the inclusion criteria defined for our review.
@@ 2. Hardware-based QoS Enforcement Features and their Origins
The rapidly increasing number of cores available on a single multi-core processor has enabled larger numbers of workloads to be co-located on a single machine. While this has its advantages, it puts increasing pressure on the resources shared among the cores of a system. Previous studies (dasari_identifying_2013; zhuravlev_addressing_2010; lofwenmark_understanding_2016) have identified the main sources of performance interference in multi-core systems. dasari_identifying_2013 (dasari_identifying_2013) identified several main sources of unpredictable delays due to resource interference. They conclude that shared main memory, shared cache, and shared interconnection networks are the main sources of unpredictability. These claims are supported by lofwenmark_understanding_2016 (lofwenmark_understanding_2016). Furthermore, zhuravlev_addressing_2010 (zhuravlev_addressing_2010) found in their (zhuravlev_addressing_2010) study that shared cache, main memory bus, and memory access controller congestion were all prominent sources of interference.
To address the problem of shared cache interference, iyer_cqos_2004 introduced the notion of QoS in the context of shared cache management in (iyer_cqos_2004) (iyer_cqos_2004). They theorized an early cache partitioning approach that was evaluated using cache simulators. While this proposal was rudimentary, it was developed upon and more refined proposals of cache and memory partitioning approaches were later published (iyer_qos_2007). Simultaneously, low-overhead shared cache monitoring solutions were also theorized (zhao_cachescouts_2007).
In more recent years, chip manufacturers have started to implement cache and memory bandwidth management features in hardware. Intel refers to these features as Intel Resource Director Technology (RDT) (intel_corporation_intel_2019), while AMD calls their implementation AMD Platform Quality of Service Extensions (QoSE) (advanced_micro_devices_inc_amd64_2022). The features allow practitioners to tune hardware to the specific needs of their workloads and thus achieve improved efficiency through reduced interference.
This section discusses the causes of performance unpredictability originating in shared resource congestion. This requires an insight into the architecture of modern multi-core processors, as well as how they implement the specific resource of interest. Additionally, the inner workings of the dynamic resource control mechanisms of modern processors are discussed.
@@ 2.1. Classes of Service
The control mechanisms of hardware-based QoS enforcement features use a concept called classes of service (COS or CLOS). This concept is grounded in hardware design and dictates how resource control rules are applied to the workloads of a system. Every logical core is assigned to one, and only one, COS at any one time. Resource control rules are set at the COS level, which will be enforced upon any core in the COS. As the system starts up, all cores are assigned to COS0, the default COS (yu_24_nodate). They may later be assigned to different COSs dynamically at runtime. The total number of available COSs is dependent on the processor model and thus changes between systems.
@@ 2.2. Shared Cache Congestion
Traditional von Neumann architecture systems have processing and memory units physically separated (arikpo2007neumann). Thus, the performance of a system becomes highly dependent on the performance of the interconnect between the two physical units, which has become known as the von Neumann bottleneck (arikpo2007neumann). Over the past few decades, advances in processing technology have increasingly outgrown that of memory technology (hennessy_computer_2011) and this difference in performance growth is referred to as the memory wall (iyer_advances_2021), which has prompted the invention of many techniques to mitigate the problem.
@@
One such technique is to use a cache. A cache is a small, high-performance memory package that is typically embedded in the CPU die. As the cache is substantially faster than the main memory, the processor may place data currently worked on in the cache and thus spend less time waiting for the data it needs from the main memory (brett_memory_2016).
@@
The caches of modern multi-core processors are so-called multi-layer caches. Each layer of cache is larger, but slower than the last. Typically, modern caches consist of three layers. Layer 1 (L1) is the smallest and fastest layer. Each physical core has a separate L1 cache split into two parts, a data cache (L1d), and an instruction cache (L1i). The next layer, layer 2 (L2), is similarly constructed so that each physical core has its own instance. However, there is no distinction between data and instruction memory in the L2 cache. Layer 3 (L3) is shared among multiple cores, often all cores on a chip. This means that there may be inter-core interference as the workloads of different cores compete for cache space. The highest numbered level of cache on a system is commonly known as the last-level cache (LLC). On a three-level cache system, LLC refers to the L3 cache. An overview of a three-layer cache architecture is illustrated in Figure 1.
Figure 1. Illustration of a typical three-layer cache architecture.
@@
Due to the cache’s small size, it is inevitable that memory will have to be read from the main memory eventually as working set sizes outgrow the available cache. AMD’s EPYC Genoa server processors with 3D V-Cache technology provide up to $1152$ megabytes of LLC cache while supporting up to 6 terabytes of main memory per socket (advanced_micro_devices_inc_amd_2023) and thus, LLC accounts for only $0.02\%$ of the main memory size. When memory required by the processor is not already present in the cache, a cache miss has occurred and a delay penalty is incurred while waiting for the required data (dasari_identifying_2013). In a (velten_memory_2022) study by velten_memory_2022, it was found that L3 access incurred a penalty of 39 cycles while waiting for main memory access took 221 cycles (velten_memory_2022)111 L3 and main memory access times are highly dependent on hardware (velten_memory_2022). The numbers presented were recorded on a system equipped with two AMD EPYC 7702 (Rome) processors. . Thus, accessing data not present in any level of cache results in a 5.67×5.67$\times$5.67 × increase in access time compared to accessing data present in the shared L3 cache.
@@
Modern data centers spend an increasing proportion of computational resources on data-intensive workloads such as machine learning (ML) training, ML inference, video processing, and audio processing. The size of the data sets used by these processes means that they are particularly destructive to cache integrity, meaning more time is spent waiting for data from the main memory. Given the shared nature of the LLC, this affects all workloads of a system and is known as shared cache congestion. Such congestion may take the form of inter-task interference when a process replaces cache lines of another process on the same core, or inter-core interference when cache lines of a process running on another core are overridden (lofwenmark_understanding_2016).
@@ 2.2.1. Shared Cache Partitioning
Configuration of the cache allocation scheme for a COS is done by setting a bitmask, the size of which is dependent on the specific CPU model. A CPU supporting $n$ bits of control may have a COS assigned $1/n$ or more [^2] COSs assigned more than one such share must on Intel platforms have all “1” bits in a contiguous block due to hardware limitations (intel_corporation_intel_2023). of the available cache. Allocation bitmasks may overlap between COSs. If so, parts of the cache are shared between two or more classes. An illustration of a cache partitioning configuration may be observed in Figure 2.
Figure 2. Illustration of a cache partitioning scheme for a CPU supporting $8$ bits of control ($0_{16}$ to $\textrm{FF}_{16}$). Here, COS0 has exclusive access to $2/8=1/4$ of available cache, while COS1 and COS2 share parts of their allocation.
When moving a core to a new COS, the new rules only apply to future cache allocations. Cache lines already present in the old partition may continue to be used until evicted (yu_24_nodate).
@@ 2.3. Shared Memory Congestion
Although techniques such as caching mitigate the von Neumann bottleneck described in Section 2.2, they do not remove the need to transfer data to and from the main memory. The interconnect between the CPU and the main memory, the memory bus, is limited in bandwidth and shared by all tasks on a non-uniform memory access (NUMA) node. Therefore, memory requests may be required to do additional waiting until sufficient capacity to handle the request is available (pellizzoni_worst_2010). This may result in workloads of low to medium memory intensity becoming drowned out given co-located workloads of high enough memory intensity. An abstract model of the interconnect between the cores, LLC, and main memory is illustrated in Figure 3.
@@
Figure 3. Illustration of the shared memory bus of a modern Intel Xeon Scalable processor (mulnix_intel_2022).
@@
Memory controllers commonly implement a variation of a first-come-first-served (FCFS) request scheduling algorithm that prioritizes row hit requests (rixner_memory_2000). This prioritized queue of requests in the memory controller influences the memory access times of processes (kim_bounding_2016).
2.3.1. Shared Memory Limiting
@@
Partitioning of the memory bus is available on supported hardware, which functionality is provided by a programmable request rate controller implemented in chip hardware. The feature is referred to as memory bandwidth allocation (MBA) by Intel (intel_corporation_intel_2023) and the Linux kernel documentation (yu_24_nodate). However, a more descriptive name is bandwidth limitation as used by AMD (advanced_micro_devices_inc_amd64_2022).
@@
Bandwidth limitation provides the ability to dynamically throttle and delay memory requests from specific processor cores. Typically, a memory-intensive process would have its memory requests limited so that other workloads become less affected by the high bandwidth usage. Bandwidth limits may be specified as a percentage (Intel only), or as a maximum bandwidth in megabytes (yu_24_nodate).
@@
Throttling is applied to all requests that miss L2 cache, the last layer of cache specific to a single physical core. This includes requests that may hit the shared L3 cache (xiang_emba_2019) as illustrated by Figure 4. Thus, bandwidth limitation is the most effective at throttling workloads that are inefficient at utilizing the L3 cache. Throttling is applied at the level of physical cores and any core utilizing simultaneous multi-threading will therefore use the maximum throttling value of its logical cores (herdrich_introduction_2019).
@@
Figure 4. Illustration of memory bandwidth limiting (MBL) impact on the memory request chain. Bandwidth limiting affects all requests that miss the L2 cache, which includes potential L3 hits.

### v2

# Hardware-Level QoS Enforcement Features: Technologies, Use Cases, and Research Challenges

@@ Abstract.
© 2025 Copyright held by the owner/author(s). This is the author accepted manuscript of a paper published in ACM Computing Surveys (CSUR), open access under the Creative Commons Attribution (CC BY 4.0) license. The final version is available at the ACM Digital Library: https://doi.org/10.1145/3774317.
@@ 1. Introduction
Support for management and partitioning of shared system resources such as shared cache and memory bandwidth has recently become available in commercial off-the-shelf x86 server processors. The features, which we refer to as hardware-based quality-of-service (QoS) enforcement features, may be dynamically controlled at runtime using processor model-specific registers or an operating system-level interface. They enable systems to enforce availability guarantees and mitigate the interference caused by shared resource congestion. The benefits of intelligently tuning such features have been shown in previous work to include increased system performance (lo_heracles_2015, 61), improved system utilization (patel_clite_2020, 76), and reduced power consumption (nejat_cooperative_2022, 68). However, to the best of our knowledge, there has been no previous attempt at surveying the academic and industrial landscapes to understand the applicability of these new features in different contexts.
@@
One such context of interest is that of general-purpose cloud computing. The cloud-connected nature of day-to-day applications implies that performance issues of cloud-based web services now have a prevalent impact on the daily lives of the general population in a way that was unthinkable a few decades ago. In fact, a 2023 US-based survey indicated that $48\%$ of companies use industry-tailored clouds with another $20\%$ intending to make the transition to cloud-based solutions within 12 months (ballard_industry_2023, 7). Latency-sensitive web services are often referred to as latency critical (LC) tasks, which is one of two broad categories cloud-based workloads may be classified into. The response times of LC services often directly affect the end-user experience and are therefore monitored and kept to performance targets specified by service level objectives (SLOs). Best-effort (BE) jobs on the other hand are time non-critical, predictable workloads that may run whenever sufficient resources become available, commonly using a batch-format scheduling system. Thus, such jobs are well suited to improve server utilization by consuming resources left unused by LC services based on current loads, a technique leveraged by several previous works (lo_heracles_2015, 61, 17, 70, 76, 73, 83, 85, 123, 57, 125, 14). BE tasks such as machine learning training or big data analysis jobs are often data intensive and consequently disproportionately “noisy” users of shared system resources. Unless mitigated, this congestion may negatively affect the performance of LC services (dasari_identifying_2013, 23).
@@
The congestion of shared resources such as last-level cache (LLC) and memory bandwidth (MB) caused by the co-location of workloads on multicore systems has been shown in previous studies to reduce system performance (dasari_identifying_2013, 23, 62, 127). The importance of this becomes evident when considering the SLOs that web services often are required to meet. Not reaching performance targets due to resource congestion may end up costly as BE jobs may need to be canceled and re-queued. Alternatively, additional computing resources may need to be provisioned if permitted by the underlying infrastructure platform. Hardware-based QoS enforcement features are designed to mitigate the issues of shared resource congestion. However, hardware-based QoS enforcement is indirect in nature (see Section 2) and thus, intelligent tuning is required to achieve predictable and performant service operations in cloud environments.
@@ 1.1. Our Contributions
    We survey a search space of $3685$ articles and identify $78$ papers that fit the inclusion criteria defined for our review.
@@
Our proposed taxonomy addresses a significant gap in the current literature. It serves as a foundation for understanding the state-of-the-art research and provides a novel overview of the applicability of hardware-based QoS enforcement features across various contexts. Additionally, we identify a lack of adoption of QoS enforcement in general-purpose cloud computing. By analyzing the literature, we aim to understand the common limitations and challenges that hinder the practicality and widespread adoption of QoS enforcement in the cloud. To this end, we present trends in the research landscape and identify challenges that need to be addressed to facilitate wider adoption of QoS enforcement in autonomous cloud computing environments.
@@ 2. Hardware-based QoS Enforcement Features and their Origins
The rapidly increasing number of cores available on a single multi-core processor has enabled larger numbers of workloads to be co-located on a single machine. While this has its advantages, it puts increasing pressure on the resources shared among the cores of a system. Previous studies (dasari_identifying_2013, 23, 127, 62) have identified the main sources of performance interference in multi-core systems. (dasari_identifying_2013, 23) (dasari_identifying_2013, 23) identified several main sources of unpredictable delays due to resource interference. They conclude that shared main memory, shared cache, and shared interconnection networks are the main sources of unpredictability. These claims are supported by (lofwenmark_understanding_2016, 62) (lofwenmark_understanding_2016, 62). Furthermore, (zhuravlev_addressing_2010, 127) (zhuravlev_addressing_2010, 127) found in their 2010 study that shared cache, main memory bus, and memory access controller congestion were all prominent sources of interference.
To address the problem of shared cache interference, (iyer_cqos_2004, 48) introduced the notion of QoS in the context of shared cache management in 2004 (iyer_cqos_2004, 48). They theorized an early cache partitioning approach that was evaluated using cache simulators. While this proposal was rudimentary, it was developed upon and refined proposals of cache and memory partitioning approaches were later published (iyer_qos_2007, 49). Simultaneously, low-overhead cache monitoring solutions were also theorized (zhao_cachescouts_2007, 124).
In more recent years, chip manufacturers have started to implement cache and memory bandwidth management features in hardware. Intel refers to these features as Intel Resource Director Technology (RDT) (intel_corporation_intel_2019, 47), while AMD calls their implementation AMD Platform Quality of Service Extensions (QoSE) (advanced_micro_devices_inc_amd64_2022, 2). The features allow practitioners to tune hardware to the specific needs of their workloads and thus achieve improved efficiency through reduced interference.
This article’s scope focuses on x86-based implementations of hardware-level QoS enforcement. However, outside the domain of x86 processors, similar features have been implemented in the ARM ecosystem by Cavium (now Marvell) (wang_swap_2017, 105) as proprietary features of their processors. More recently, official ARM support for QoS enforcement features has been made available through the Memory System Resource Partitioning and Monitoring (MPAM) technology (arm_limited_arm_2024, 4). Additionally, efforts have been made to introduce similar features to RISC-V platforms (zhang_labeled_2022, 120).
The remainder of this section discusses the causes of performance unpredictability originating in shared resource congestion. This requires an insight into the architecture of modern multi-core processors, as well as how they implement the specific resource of interest. Additionally, the inner workings of the dynamic resource control mechanisms of modern x86 processors are discussed.
@@ 2.1. Classes of Service
The control mechanisms of hardware-based QoS enforcement features use a concept called classes of service (COS or CLOS). This concept is grounded in hardware design and dictates how resource control rules are applied to the workloads of a system. Every logical core is assigned to one, and only one, COS at any one time. Resource control rules are set at the COS level, which will be enforced upon any core in the COS. As the system starts up, all cores are assigned to COS0, the default COS (yu_24_nodate, 116). They may later be assigned to different COSs dynamically at runtime. The total number of available COSs is dependent on the processor model and thus changes between systems.
@@ 2.2. Shared Cache Congestion
Traditional von Neumann architecture systems have processing and memory units physically separated (arikpo2007neumann, 3). Thus, the performance of a system becomes highly dependent on the performance of the interconnect between the two physical units, which has become known as the von Neumann bottleneck (arikpo2007neumann, 3). Over the past few decades, advances in processing technology have increasingly outgrown that of memory technology (hennessy_computer_2011, 40) and this difference in performance growth is referred to as the memory wall (wulf_hitting_1995, 107), which has prompted the invention of many mitigating techniques.
@@
One such technique is to use a cache. A cache is a small, high-performance memory package that is typically embedded in the CPU die. As the cache is substantially faster than the main memory, the processor may place data currently worked on in the cache and thus spend less time waiting for the data it needs from the main memory (brett_memory_2016, 10).
@@
The caches of modern multi-core processors are so-called multi-level caches. Each level of cache is larger, but slower than the last. Typically, modern caches consist of three levels. Level 1 (L1) is the smallest and fastest level. Each physical core has a separate L1 cache split into two parts, a data cache (L1d), and an instruction cache (L1i). The next level, level 2 (L2), is similarly constructed so that each physical core has its own instance. However, there is no distinction between data and instruction memory in the L2 cache. Level 3 (L3) is shared among multiple cores, often all cores on a chip. This means that there may be inter-core interference as the workloads of different cores compete for cache space. The highest numbered level of cache on a system is commonly known as the last-level cache (LLC). On a three-level cache system, LLC refers to the L3 cache. An overview of a three-level cache architecture is illustrated in Figure 1.
Figure 1. Illustration of a typical three-level cache architecture.:
@@
Due to the cache’s small size, it is inevitable that memory will have to be read from the main memory eventually as working set sizes outgrow the available cache. AMD’s EPYC Genoa server processors with 3D V-Cache technology provide up to $1152$ megabytes of LLC cache while supporting up to 6 terabytes of main memory per socket (advanced_micro_devices_inc_amd_2023, 1) and thus, LLC accounts for only $0.02\%$ of the main memory size. When memory required by the processor is not already present in the cache, a cache miss has occurred and a delay penalty is incurred while waiting for the required data (dasari_identifying_2013, 23). In a 2022 study by (velten_memory_2022, 104), it was found that L3 access incurred a penalty of 39 cycles while waiting for main memory access took 221 cycles (velten_memory_2022, 104) [^1] L3 and main memory access times are highly dependent on hardware (velten_memory_2022, 104). The numbers presented were recorded on a system equipped with two AMD EPYC 7702 (Rome) processors. . Thus, accessing data not present in any level of cache results in a $5.67\times$ increase in access time compared to accessing data present in the shared L3 cache.
@@
Modern data centers spend an increasing proportion of computational resources on data-intensive workloads such as machine learning (ML) training, ML inference, video processing, and audio processing. The size of the data sets used by these processes means that they are particularly destructive to cache integrity, meaning more time is spent waiting for data from the main memory. Given the shared nature of the LLC, this affects all workloads of a system and is known as shared cache congestion. Such congestion may take the form of inter-task interference when a process replaces cache lines of another process on the same core, or inter-core interference when cache lines of a process running on another core are overridden (lofwenmark_understanding_2016, 62).
@@ 2.2.1. Shared Cache Partitioning
Configuration of the cache allocation scheme for a COS is done by setting a bitmask, the size of which is dependent on the specific CPU model. Typically, each bit corresponds to a separate hardware cache way (sohal_closer_2022, 95), but this is not necessarily the case (yu_24_nodate, 116). A CPU supporting $n$ bits of control may have a COS assigned $1/n$ or more [^2] COSs assigned more than one such share must on Intel platforms have all “1” bits in a contiguous block due to hardware limitations (intel_corporation_intel_2023, 46). of the available cache. Allocation bitmasks may overlap between COSs. If so, parts of the cache are shared between two or more classes. An illustration of a cache partitioning configuration may be observed in Figure 2.
Figure 2. Illustration of a cache partitioning scheme for a CPU supporting $8$ bits of control ($0_{16}$ to $\textrm{FF}_{16}$). Here, COS0 has exclusive access to $2/8=1/4$ of available cache, while COS1 and COS2 share parts of their allocation.:
When moving a core to a new COS, the new rules only apply to future cache allocations. Cache lines already present in the old partition may continue to be used until evicted (yu_24_nodate, 116).
@@ 2.3. Shared Memory Congestion
Although techniques such as caching mitigate the von Neumann bottleneck described in Section 2.2, they do not remove the need to transfer data to and from the main memory. The interconnect between the CPU and the main memory, the memory bus, is limited in bandwidth and shared by all tasks on a non-uniform memory access (NUMA) node. Therefore, memory requests may be required to do additional waiting until sufficient capacity to handle the request is available (pellizzoni_worst_2010, 77). This may result in workloads of low to medium memory intensity becoming drowned out given co-located workloads of high enough memory intensity. An abstract model of the interconnect between the cores, LLC, and main memory is illustrated in Figure 3.
@@
Figure 3. Illustration of the shared memory bus of a modern Intel Xeon Scalable processor (mulnix_intel_2022, 66).:
@@
Memory controllers commonly implement a variation of a first-come-first-served (FCFS) request scheduling algorithm that prioritizes row hit requests (rixner_memory_2000, 88). This prioritized queue of requests in the memory controller influences the memory access times of processes (kim_bounding_2016, 53).
2.3.1. Shared Memory Bandwidth Limiting
@@
Partitioning of the memory bus is available on supported hardware, which functionality is provided by a programmable request rate controller implemented in chip hardware. The feature is referred to as memory bandwidth allocation (MBA) by Intel (intel_corporation_intel_2023, 46) and the Linux kernel documentation (yu_24_nodate, 116). However, a more descriptive name is bandwidth limitation as used by AMD (advanced_micro_devices_inc_amd64_2022, 2).
@@
Bandwidth limitation provides the ability to dynamically throttle and delay memory requests from specific processor cores. Typically, a memory-intensive process would have its memory requests limited so that other workloads become less affected by the high bandwidth usage. Bandwidth limits may be specified as a percentage (Intel only), or as a maximum bandwidth in megabytes (yu_24_nodate, 116).
@@
Throttling is applied to all requests that miss L2 cache, the last level of cache specific to a single physical core. This includes requests that may hit the shared L3 cache (xiang_emba_2019, 109) as illustrated by Figure 4. Thus, bandwidth limitation is the most effective at throttling workloads that are inefficient at utilizing the L3 cache. Throttling is applied at the level of physical cores and any core utilizing simultaneous multi-threading will therefore use the maximum throttling value of its logical cores (herdrich_introduction_2019, 42).
@@
Figure 4. Illustration of memory bandwidth limiting (MBL) impact on the memory request chain. Bandwidth limiting affects all requests that miss the L2 cache, which includes potential L3 hits.:
2.4. Indirectness of Hardware-Level QoS Enforcement
Many of the challenges of adopting hardware-level QoS enforcement features stem from the indirect nature of their design. Both cache partitioning and memory bandwidth limiting (MBL) exhibit this indirectness. Let us first consider MBL. Any given workload will have certain memory bandwidth requirements that among other things is dependent on other system workloads through congestion in shared cache. It is not possible to give direct priority to a workload to satisfy these requirements using MBL. Rather, other system workloads must be throttled to indirectly free up bandwidth for the workload in focus.
In the case of cache partitioning, the problem of indirectness is less obvious. Intuition may tell us that assigning a larger cache allocation to a workload will improve its performance. However, as was shown by (chung_enforcing_2019, 20) (chung_enforcing_2019, 20), more LLC is not always better. They found that the performance of a workload may be negatively impacted by an increased cache allocation. This is due to the fact that other co-located workloads now share a smaller portion of cache and may subsequently experience more cache misses. This increase in cache misses result in increased congestion over the memory bus, saturating its bandwidth. Thus, when the workload in focus eventually requires data from the main memory, it may be delayed due to the congestion caused by other workloads which results in a net negative performance. It is thus not possible to make assumptions about workload cache allocation without considering the indirect behavior of other co-located workloads.